What should be tension applied to a wire of length 1m and mass 10g if it has to vibrate with fundamental frequency of 50 Hz

Grade:12

2 Answers

Arun
6 years ago
Dear Akshay
 
mass m = 10 gm = 0.01
n = (1/2L) \sqrt(T/m)
Now
50 = 1/2 \sqrt(T/0.01)
100² = T/0.01
10000*0.01 = T
T = 100 N
 
Regards
Arun (askIITiana forum expert)
Phanindra
6 years ago
n = 1/2l \sqrt(T/\mu )...
Here both T and μ are under square root...
T = tension in the string (newton)
μ = linear mass density = mass/length..
      ...(kg per meter)
 l = length of the string...(in meters) 
Now substitute the values to get Tension in the string..